# What Makes a Visualization Complex?

### v1

@@ I.2.4 Object-based Measure
Since visualizations are not typically treated as object-based scenes, where elements are explicitly counted or statistical attributes are quantified, we explore the use of color to differentiate visual elements (Item (12) O.MeC, Table 2). Results from our workshop further suggest that color is a leading factor influencing image clarity. Together, these insights allow us to investigate how high-level visualization elements contribute to perceived VC.
@@ II A Crowd-Sourcing Study
Having established an understanding of the factors influencing VC through extensive literature review, we proceeded to directly collect crowdsourced VC scores. We began with a workshop and pilot studies to inform the proper procedures for our data collection.
@@ II.1 Image Data Collection and Characteristics
We prioritize diverse stimuli and therefore used both VIS30K [19] and MASSVIS [11] as primary data sources. They both come from real-world applications thus let us examine practical VC. Most importantly, they function differently: VIS30K contains scholarly images texting formal IEEE publications, often presents scientific discoveries or new design, where annotation is rare and the corresponding description often appears in the captions or main text. In contrast, MASSVIS images are used mainly for communication, therefore key insights and trends are sometimes annotated, which further serve our goal of understanding non-graphical elements, e. g., text in images. Also, VIS30K contains glyphs and vector field images not found in MASSVIS.
II.2 A Workshop and A Pilot Study
@@
The goal of the workshop was mainly to choose a reliable method to collect a large number of VC scores. Extensive detail of the activities and outcomes are provided in the Apdx. subsection C.2. Here we give a high-level overview and decisions influenced by these activities.
@@
We first adapted a recent and well-accepted method used in BeauVis for a direct score assignment [47] and found it difficult to align scores above several 10s of images (He et al. exp used 15 images). Next we conducted an in-person exploratory workshop with 49 participants, akin to the hierarchical division method of Oliva et al. [99]. Participants ranked images and captured factors influencing VC. We found that it was well-suited for collecting comments and factors, however, it also did not scale (Oliva et al. [99] used 100 images). We also found that not a single participant used a single criterion in their VC assignment.
@@
The third method we tried was an active sampling solutions that could provide us absolute VC scores via pair-wise 2AFC ranking, successfully used in the computer vision community to collect a large set of image VC scores (e. g., [124, 97]). The idea was similar to how game competitions assigned competing players to achieve a global score for each player. Our piloting experiments with 90 and 200 images supported that the single-stage Bradley-Terry ranking [13] and Microsoft’s TrueSkill™ [52] were both reliable. However, scaling up this approach to 1,000 using only one-time sampling would require many more trials, not very feasible. We finally piloted Mikhailiuk et al. [90]’s multi-stage active sampling algorithm (Apdx. Figure 10), which could achieve the same level of reliability as the single-stage active sampling with of total trials. In each phase, the algorithm optimized the information obtained from all pairwise comparisons to compute reliable VC scores, while limiting the total number of comparisons. Our pilot testing using visualization images obtained the same reliability as the authors described and saved running cost by 9 folds.
@@
In the pilot study, we also discussed what not to collect. For example, we excluded other high-level subjective metrics, such as aesthetics, understandability, and readability, though they might correlate with VC [116, 112, 139] because these factors are also inherently subjective, potentially relying on the same set of objective metrics. Instead, we focused on objective image metrics, aiming to develop an understanding that remains independent of individuals’ backgrounds, experience and expertise, for quantifying the crowdsourcing perceived VC scores. Images were static free of measuring viewers’ ability to navigate.
II.3 Participants and Procedure
@@
At this stage, we had clear ideas on the image data and the method to collect the perceived VC scores. Now we report our crowd-sourcing experiment via the online experiment platform Prolific [103] to capture the absolute perceived VC scores for the 1,800 images.
@@
We instructed participants on the experiment procedure and provided them the high-level definition of VC always shown on the screen “the amount of detail or intricacy in images” (Apdx. Figure 9) as used in Purchase et al. [107, 129]. They were instructed not to infer knowledge and not to use information other than those presented in the images, following the instruction of Oliva et al [63]. We also had two attention test trials to flag inattentive participants out of the study. In each stage, about 20-25 volunteers compared 79 pairs of images for us to determine the VC for that stage, viewed images recommended by the active sampling algorithm. Participants viewing image pairs recommended by the algorithm. At each stage, an absolute complexity score for each image was updated from a probabilistic process to generate a global complexity score described in more detail next.
@@
Participants’ browser size and their operating system were automatically detected and their screen view must be at least on a desktop operating system to be eligible. Before this data collection, we first collected participants’ demographic information and ensured accountability by recruiting only those with a success rate above on Prolific. Additionally in the post-questionnaire, we gathered participants’ written reasoning from two randomly selected trials.
@@
In total, 349 participants (aged 18 to 25 (111), 26 to 35 (143), 36 to 45 (51), 46 to 55 (27) and over 55 (17)), with education levels: graduate school (134), college (130), professional school (55) and some high school (30). The gender distribution is nearly balanced: 168 women, 176 men, and 5 who did not disclose their gender). The participants were compensated € 9.5 / hour for their participation.
@@ II.4 Crowdsourcing Results
The experiment was conducted over 14 stages for approximately 34 hours, resulting in a total of 27,571 trials. Each trial lasted an average of 4.34 seconds ( CI: [4.17, 4.50]). Each visualization image was viewed approximately 234 times ( CI: [202, 266]).
VC Score Calculation To derive VC scores for all 1,800 images from the on average 1,969 pairwise comparison data per run, the active learning first constructed a comparison matrix containing the outcomes of all pairwise comparisons, where each entry recorded how often an image was selected as more visually complex when compared against others, and then applied the TrueSkill algorithm [52], to infer relative complexity from the comparisons.
TrueSkill is a Bayesian ranking algorithm that models each image’s visual complexity as a Gaussian distribution , where represents the estimated VC score and reflects the uncertainty. During each comparison between two images and , TrueSkill estimates the probability that is more visually complex than by comparing their distributions. The outcome of the comparison updates both distributions using Bayesian inference. Specifically, the posterior means and variances are updated based on the observed result and the prior uncertainties. If image is judged more complex than , the mean is increased and is decreased; both variances and are reduced as the model gains confidence. The amount of adjustment depends on the relative uncertainties of the two images. The probability that image is more complex than image is calculated using comparison noise and a cumulative distribution function. After many pairwise comparisons, the distributions converge.
For our analyses, the final VC score of each image is taken as the posterior mean which reflects the inferred likelihood that it would be judged more complex than a randomly chosen image from the dataset. (Additional validation in Apdx. Figure 10.) Figure 2 shows a subset of 84 images, sorted in ascending order of their VC scores from left to right, then top to bottom. (Apdx. Figure 18 provides the full set of 1,800 VC images and scores. )
@@ III.1 Our Two Object-based Image Quality Metrics
We define two metrics related to the semantic information of a scene, building on the concept of graphical elements, where objects are recognized quickly, sometimes in a single glance [101]. While our community has traditionally relied on edges to define boundaries, we focus here on high-level perception of using colors, drawing on insights from vision science, including the role of color in delineating element boundaries and the preattentive nature of coloring in visualizations [21]. Also, we saw that many images contain categorical colors in the legend. These perspectives informed our decision to use color as a means of defining boundaries between data elements, alongside text, which can be recognized and quantified with relative ease.
Text-ink ratio (O.TiR) represents the proportion of text-area pixels in an image. Here, “text” refers to contextual information, including annotations, axis labels, or titles. Any text used to explain, clarify, or define image content is considered “annotated” [109]. Text-based visualizations where text is used to show data are excluded. There are many methods to compute O.TiR. In our case, we used an OCR (Optical Character Recognition) package PP-OCR [34] to compute text regions and then use the VIS30K bounding box labeling tool [19] (Apdx. Figure 21) to curate annotations free of algorithmic errors.
Color count (O.MeC) defines the number of distinct colors/objects in a visualization image. Several considerations guided how we computed the MeC, due to the difficulty to name the groups as Rosenholtze et al. pointed out [120]. For example, would a data sample in a scatterplot be considered an item or a cluster of these an ‘object’. Here, we assume that designers would have distinguished ‘objects’ by color—if the goal is to see the cluster, they would be shown as color-based clusters. As a result, for discrete data, we simply count the number of object by looking up legend without concerning items within that cluster, since legend captures human attention first [14] and color supports fast pop-up processing [147, 158]. For these images using discrete colors, we count the number of unique colors present in the legend, along with any additional text color.
@@
One may argue that luminance areas were not consistently counted in our analysis. In practice, we addressed this on a case-by-case basis. For texture-based pattern images such as LIC, we included luminance in the color count. Since the H-S naming system is not uniformly distributed in the L*A*B space, some color names are perceptually closer than others. In cases like this, we applied a color distance threshold: colors with a E were merged (Apdx. Figure 22). This value is slightly higher than the typical human perceptual threshold ( 10 [152]), allowing us to account for pixel-level noise and small color variations. This threshold also helped reduce salt-and-pepper noise in the image. Two co-authors curated the final answers.
@@ III.2 Observing 12 Objective Image Quality Metrics
Information-theoretic-based metrics. We used four metrics: (1) the original Shannon Image Entropy (O.IE) [127]; (2) Kolmogorov Complexity (O.KC), which estimates an upper bound of an image’s information capacity using the zstd compression method [25], where higher complexity indicates greater internal redundancy, resulting in larger compressed sizes; (3) Subband entropy (O.SE) [118], measuring the randomness of information within different frequency components of an image after wavelet decomposition; and (4) information gain (O.IG) [3], which examines neighboring pixel probabilities and sensitive to local textures, edges, and details within the image. Here, We observed that these four information-theoretic metrics are highly correlated emphasizing similar regions of an image (Figure 4 and Apdx. Figure 13).
Clutter-based Metrics include (5) feature congestion (O.FC) [118] and (6) heterogeneity (O.H) [29], considering arrangement and overall pattern distribution. O.FC quantifies how "cluttered" or "busy" an image appears by analyzing structural variations of color and luminance contrast. Similarly, higher O.H indicate greater contrast and variation in gray levels, which suggest a more complex or textured image. Figure 4 shows that areas with high density pattern and structure variations were emphasized by these two metrics in the images below.
Color-based metrics. For O.ERGB, higher entropy indicates greater color variation. We can see that smaller regions with many color changes tend to have higher ERGB, capturing the edges between the colored bars. In contrast, grey-scale images exhibit lower scores due to reduced color variability. O.ERGB also successfully highlighted nodes in node-link diagrams. On the other hand, O.CF tends to miss dense node areas but causes empty regions to stand out more prominently.
Shape-based metrics. We used (9) edge density (O.ED) studied in [120, 86, 118], representing the ratio of edge pixels with discontinuities in image intensity to total pixels in an image and (10) feature points (O.FP) [44] represent the corner points with significant variation from surrounding pixels. They both emphasize areas with highly dense lines and significant variations in color values.
Multi-colinearity analysis. We also performed the co-linearity analysis (see Apdx. Figure 13 for detailed analysis). The main take away message was that we observed within category correlation especially in the pixel-based information theoretic and between shape-based metrics. In contrast, the object-based metrics, O.TiR and O.MeC showed little to no correlation with other metrics.
@@ III.3.1 Analysis Method
We computed PLS regression coefficients to estimate the relative contributions of each metric to perceived VC. We also used bootstrap resampling to estimate the variability of each coefficient and compute -values, thereby identifying which metrics are more significant and stable across resamples (details in Apdx. subsection C.3). We used the resulting PLS regression coefficients to estimate the relative contribution of each metric to perceived VC. Interpreting these coefficients, however, requires care. Unlike in standard regression, the coefficients in PLS reflect the contribution of a variable conditional on its shared variance with other predictors, which means the following: (1) variables with larger absolute coefficient values are interpreted as having greater influence, while the sign of the coefficient indicated whether the effect was positive or negative; (2) small coefficients may not indicate that a metric is unimportant, but rather that its influence overlaps with other variables or too unstable to predict VC.
@@ III.3.2 Results: Overall Factorizing Perceived VC
(2) Information-theoretic metrics based on specific stimulus and attributes are significant factors. Specifically, O.FC in the clutter category, adapted to statistical distributions of color and texture, and O.ERGB in the color category, the original Shannon entropy on color alone, were both significant main effects with () and (), respectively. The other metric in the clutter category, O.H, that computed pixel-level heterogeneity of an image, was also significant due to a small variation in data, although its coefficient was small (). The metric that measured the variety and diversity of colors (O.CF), however, did not show a significant contribution to perceived VC prediction.
@@
(4) Text-ink-ratio reduces perceived visual complexity. O.TiR, which measured the proportion of annotated text within an image, was found to have a negative coefficient (-0.12, p0.005), suggesting that the presence of annotation may in general reduce rather than increase perceived VC. As discussed earlier, annotated text can enhance clarity and aid in interpretation, thereby lowering perceived complexity. To explore this further, we analyzed the trend between O.TiR and perceived VC and observed a non-linear, bell-shaped relationship (Figure 6). Specifically, VC initially decreases as O.TiR increases, reaching a minimum around the range, and rises again beyond a threshold of approximately . This difference in complexity is statistically significant (), indicating a potential optimal range of O.TiR, where a moderate amount of text can reduce perceived complexity by improving clarity, while too little or too much text may have the opposite effect.
We also observed that increasing O.MeC was associated with higher perceived VC (Figure 6). Among the dataset, we found 96 black-and-white or monochrome images with no color variation (). While natural scenes are almost always colorful, visualization images can achieve similar perceptual differentiation through other cues such as texture and shape variation. In general, grayscale or low-color images were perceived as less complex than colored ones, a difference that was statistically significant (). For the images with higher O.MeC, we found that color often served as a boundary-defining cue, helping to visually separate elements. This boundary-forming function of color contributed to increased VC, especially when the colors clearly delineated objects or categories. Notably, this effect was consistent regardless of whether the image included textual annotations ( for both cases). Among these color-rich images, some used discrete colormaps with clearly defined edges, while others employed continuous colormaps such as gradient-style heatmaps. In some cases, these color patterns also reinforced coherent textural structures, further adding to visual complexity. Due to these observations, we next separate color/texture images by continuity to further investigate their relationship with perceived VC.
@@
Our analysis shows that all variable groups contribute to higher perceived visual complexity: information-theoretic metrics, greater edge density, more feature points, and a larger number of distinct objects. This result replicates the findings in vision science that visual complexity is likely to be a high-dimensional phenomenon influenced by many factors. As shown in the three panels of What Makes a Visualization Complex?, the simplest and most complex visualizations differ notably in visual content. Similar patterns are observed in Apdx. Figure 17 when comparing samples from MASSVIS and VIS30K. Some visualizations, particularly Matrix views, appear more complex due to their use of many small points and a wider range of colors. In contrast, simpler images tend to contain fewer elements and use minimal colors.

### v3

@@ I.2.4 Object-based Measures
Since images are not typically treated as object-based scenes, where elements are explicitly counted or statistical attributes are quantified, we explore the use of color to differentiate visual elements (Item (12) O.MeC, Table 2). Results from our workshop further suggest that color is a leading factor influencing image clarity. Together, these insights allow us to investigate how high-level visualization elements contribute to perceived VC.
@@ II A Crowd-Sourcing Study
Having established an understanding of the factors influencing VC through an extensive literature review, we proceeded to directly collect crowdsourced VC scores. We began with a workshop and pilot studies to inform the proper procedure for our data collection.
@@ II.1 Image Data Collection and Characteristics
We prioritize diverse stimuli and therefore used both VIS30K [18] and MASSVIS [10] as primary data sources. They both come from real-world applications thus let us examine practical VC. Most importantly, they function differently: VIS30K contains scholarly images extracted from formal IEEE publications, and often presents scientific discoveries or new designs, where annotation is rare and the corresponding description often appears in the captions or main text. In contrast, MASSVIS images are used mainly for communication, therefore key trends are sometimes annotated, which further serve our goal of understanding non-graphical elements, e. g., text in images. Also, VIS30K contains glyphs and vector field images not found in MASSVIS.
II.2 Workshop and Pilot Study
@@
The goal of the workshop was mainly to choose a reliable method to collect a large number of VC scores. Extensive details of the activities and outcomes are provided in Apdx. subsection C.2. Here we give a high-level overview and decisions influenced by these activities.
@@
We first adapted a recent method used in BeauVis for a direct score assignment [45], but found it difficult to align scores above several 10s of images (He et al. [45] used 15 images). Next, we conducted an in-person exploratory workshop with 49 participants, akin to the hierarchical division method of Oliva et al. [98]. Participants ranked images and captured factors influencing VC. We found that it was well-suited for collecting comments and factors, however, it also did not scale (Oliva et al. [98] used 100 images). We also found that not a single participant used a single criterion in their VC assignment.
@@
The third method we tried was an active sampling solution that could provide us absolute VC scores via pairwise 2AFC ranking successfully used in the computer vision community to collect a large set of image VC scores (e. g., [124, 96]). The idea was similar to how game competitions assigned competing players to achieve a global score for each player. Our piloting experiments with 90 and 200 images supported that the single-stage Bradley-Terry ranking [12] and Microsoft’s TrueSkill™ [50] were both reliable. However, scaling up this approach to 1,000s in the global space using only one-time sampling would require many more trials, which was not feasible. We finally piloted Mikhailiuk et al. [89]’s multi-stage active sampling algorithm (Apdx. Figure 10), which could achieve the same level of reliability as the single-stage active sampling with of total trials. In each phase, the algorithm optimized the information obtained from all pairwise comparisons to compute reliable VC scores, while limiting the total number of comparisons. Our pilot tests achieved the same reliability as the authors reported while reducing operating costs by 9 fold.
@@
In the pilot study, we also discussed what not to collect. For example, we excluded other high-level subjective metrics, such as aesthetics, understandability, and readability, even though they might correlate with VC [116, 112, 141] because these factors are also inherently subjective, potentially relying on the same set of objective metrics. Instead, we focused on objective image metrics, aiming to develop an understanding that remains independent of individuals’ experience and expertise, for quantifying the crowdsourced perceived VC scores. Images were static and not dependent on viewers’ ability to navigate.
II.3 Participants and Multi-stage Procedure
@@
At this stage, we had clear ideas on the image data and the method to collect the perceived VC scores. Now we report our crowd-sourcing experiment via the online experiment platform Prolific [102] to capture the absolute perceived VC scores.
@@
We instructed participants on the experiment procedure and provided them the high-level definition of VC, “the amount of detail or intricacy in images” (Apdx. Figure 9) as used in Purchase et al. [107, 130], which was always displayed on the screen. They were instructed not to infer knowledge and not to use information other than those presented in the images, following the instruction of Oliva et al [61]. We also had two attention test trials to identify and remove inattentive participants. In each stage, about 20-25 volunteers compared 79 pairs of images to determine the VC for that stage, viewing images recommended by the active sampling algorithm. At each stage, an absolute complexity score for each image was updated from a probabilistic process to generate a global complexity score described in more detail next.
@@
Participants’ browser type and operating system were automatically detected, and the screen resolution must be at least on a desktop operating system. Before this data collection, we first collected participants’ demographic information and ensured accountability by recruiting only those with a success rate above on Prolific. Additionally in the post-questionnaire, we gathered participants’ written reasoning from two randomly selected trials.
@@
In total, there were 349 participants with ages distributed as follows: 18-25 (111), 26-35 (143), 36-45 (51), 46-55 (27), and over 55 (17). Their education levels included graduate school (134), college (130), professional school (55), and some high school (30). The gender distribution was nearly balanced, with 168 women, 176 men, and 5 who did not disclose their gender. The participants were compensated € 9.5 / hour for their participation.
@@ II.4 Crowdsourcing Results
The experiment was conducted over 14 stages for approximately 34 hours, resulting in a total of 27,571 trials, with an average of 1,969 trials per stage. Each trial lasted an average of 4.34 seconds ( CI: [4.17, 4.50]). Each visualization image was viewed approximately 234 times ( CI: [202, 266]).
VC Score Calculation. We applied the Mikhailiuk et al.’s approach [89] to calculate the VC scores for all 1,800 images. The approach first resulted in a comparison matrix containing the outcomes of all pairwise comparisons, where each entry recorded how often an image was selected as more visually complex when compared against others, and then applied the TrueSkill™ algorithm [50], to infer relative complexity from the comparisons. TrueSkill™ is a Bayesian ranking algorithm that models each image’s VC as a Gaussian distribution with a mean () representing the estimated VC score and a standard deviation () representing the uncertainty of that estimate in our context. During each comparison between two images and , TrueSkill™ estimates the probability that is more visually complex than . Specifically, if image is judged more complex than , the mean is increased and is decreased; both variances and are reduced as the model gains confidence through increased comparisons. After many pairwise comparisons, the distributions converge.
For our analysis, the final VC score of each image is taken as the posterior mean which reflects the inferred likelihood that it would be judged more complex than a randomly chosen image from the dataset (Additional validation in Apdx. Figure 10). Figure 2 shows a subset of 84 images, sorted in ascending order of their VC scores from left to right, then top to bottom (Apdx. Figure 11 provides more images and scores).
@@ III.1 Our Two Object-based Image Quality Metrics
We define two metrics related to the semantic information of a scene, building on the concept of graphical elements, where objects are recognized quickly, sometimes in a single glance [100]. While our community has traditionally relied on edges to define boundaries, we focus here on high-level perception of using colors. Here, we drew on insights from vision science, including the role of color in delineating element boundaries and the preattentive nature of coloring in visualizations [21]. Also, we saw that many images contained categorical colors in the legend. These perspectives informed our decision to use color as a means of defining boundaries between data elements, alongside text, which can be recognized and quantified with relative ease.
Text-ink ratio (O.TiR) represents the proportion of text-area pixels in an image. Here, “text” refers to contextual information, including annotations, axis labels, or titles. Any text used to explain, clarify, or define image content is considered “annotated” [109]. We excluded text-based visualizations where text was used to show data. There are many methods to compute O.TiR. In our case, we used an OCR (Optical Character Recognition) package PP-OCR [77] to compute text regions and then use the VIS30K bounding box labeling tool [18] (Apdx. Figure 13) to curate annotations, legends, etc.
Color count (O.MeC) defines the number of distinct colors/objects in a visualization image. Several considerations guided how we computed the MeC, due to the difficulty to name the groups as Rosenholtz et al. pointed out [120]. For example, would a data sample in a scatterplot be considered an item or a cluster of these an item? Here, we assume that designers would have distinguished ‘items’ by color—if the goal is to see the cluster, they would be shown as color-based clusters. As a result, for discrete data, we simply count the number of items by looking at the legend without concerning items within that cluster, since legend captures human attention first [13] and color supports fast pre-attentive processing [149, 160]. For these images using discrete colors, we count the number of unique colors present in the legend, along with any additional text color.
@@
One may argue that luminance areas were not consistently counted in our analysis. In practice, we addressed this on a case-by-case basis. For texture-based pattern images such as LIC, we included luminance in the color count. Since the H-S naming system is not uniformly distributed in the L*A*B space, some color names are perceptually closer than others. In cases like this, we applied a color distance threshold: colors with a E were merged (Apdx. Figure 23). This value is slightly higher than the typical human perceptual threshold ( [154]), allowing us to account for pixel-level noise and small color variations. This threshold also helped reduce salt-and-pepper noise in the image. Two ( and last) co-authors curated the final answers.
@@ III.2 Observing 12 Objective Image Quality Metrics
We applied the 12 metrics to our dataset and had a few observations.
Information-theoretic-based metrics. We used six metrics. Among these four were the original Shannon’s entropy or pixel-based framed as: (1) the original Shannon Image Entropy (O.IE) [128]; (2) Kolmogorov Complexity (O.KC), which estimates an upper bound of an image’s information capacity using the zstd compression method [25], where higher complexity indicates greater internal redundancy, resulting in larger compressed sizes; (3) Subband entropy (O.SE) [119], measuring the randomness of information within different frequency components of an image after wavelet decomposition; and (4) information gain (O.IG) [3], which examines neighboring pixel probabilities and is sensitive to local textures, edges, and details within the image. We observed that these four information-theoretic metrics were highly correlated emphasizing similar regions of an image (Figure 4 and Apdx. Figure 22).
Clutter-based Metrics include a statistic information theoretic metric: (5) feature congestion (O.FC) [119]. O.FC quantifies how busy an image appears by analyzing structural variations of color and luminance contrast using information-theoretic measure. In contrast, (6) heterogeneity (O.H) [29] considers arrangement and overall pattern distribution, and higher O.H indicates greater contrast and variation in gray levels, which suggest a more complex or textured image. Figure 4 shows that areas with a high density pattern and structure variations were emphasized by these two metrics.
Color-based metrics. O.ERGB is the last information-theoretic metric, and higher entropy indicates greater color variation. We can see that smaller regions with many color changes tend to have higher ERGB, capturing the edges between the colored bars. O.ERGB also highlighted nodes in node-link diagrams. On the other hand, colorfulness (O.CF) tends to make the larger areas stand out more prominently.
Shape-based metrics. We used (9) edge density (O.ED), as studied in [120, 85, 119], which represents the ratio of edge pixels—indicating discontinuities in image intensity—to total number of pixels in an image, and (10) feature points (O.FP) [42], which are corner points exhibiting significant variation from surrounding pixels. Both metrics emphasize areas with densely packed lines and significant variations in color values.
Multi-colinearity analysis. We also performed a co-linearity analysis (see Apdx. Figure 22 for detailed analysis). We observed correlations within categories, especially among pixel-based information-theoretic metrics, and between shape-based metrics. In contrast, the object-based metrics, O.TiR and O.MeC showed little to no correlation with other metrics.
@@ III.3.1 Analysis Method
We computed PLS regression coefficients to estimate the relative contributions of each metric to perceived VC. We also used bootstrap resampling to estimate the variability of each coefficient and compute -values, thereby identifying which metrics are more significant and stable across resamples (details in Apdx. subsection C.3). We used the resulting PLS regression coefficients to estimate the relative contribution of each metric to perceived VC. Interpreting these coefficients, however, requires care. Unlike in standard regression, the coefficients in PLS reflect the contribution of a variable conditional on its shared variance with other predictors, which means the following: (1) variables with larger absolute coefficient values are interpreted as having greater influence, while the sign of the coefficient indicates whether the effect is positive or negative; (2) small coefficients may not indicate that a metric is unimportant, but rather that its influence overlaps with other variables or is too unstable to predict VC.
@@ III.3.2 Results: Overall Factorizing Perceived VC
(2) Information-theoretic metrics based on specific stimulus and attributes are significant factors. The pixel-based information-theoretic metrics (O.IE, O.KC, and O.IG) were not the significant main effect, but O.SE using the frequency was. In addition, O.FC in the clutter category and O.ERGB in the color category are also information-theoretic metrics. O.FC adapted subband entropy to statistical distributions of color and texture and O.ERGB uses the original Shannon entropy on color alone. They were both significant main effects with () and (), respectively.
@@
(4) Text-ink-ratio reduces perceived visual complexity. O.TiR, which measured the proportion of annotated text within an image, was found to have a negative coefficient (-0.12, p0.005), suggesting that the presence of annotation may in general reduce rather than increase perceived VC. As discussed earlier, annotated text can enhance clarity and aid in interpretation, thereby lowering perceived complexity. To explore this further, we analyzed the trend between O.TiR and perceived VC and observed a non-linear, bell-shaped relationship (Figure 6). Specifically, VC initially decreases as O.TiR increases, reaching a minimum , and rises again beyond a threshold of approximately . This difference in complexity is statistically significant (), indicating a potential optimal range of O.TiR, where a moderate amount of text can reduce perceived complexity by improving clarity, while too little or too much text may have the opposite effect.
We also observed that increasing O.MeC was associated with higher perceived VC (Figure 6). Among the dataset, we found 96 black-and-white or monochrome images with no color variation (). While natural scenes are almost always colorful, visualization images can achieve similar perceptual differentiation through other cues such as texture and shape variation. In general, grayscale or low-color images were perceived as less complex than colored ones, a difference that was statistically significant (). For the images with higher O.MeC, we found that color often served as a boundary-defining cue, helping to visually separate elements. This boundary-forming function of color contributed to increased perceived complexity, especially when the colors clearly delineated objects or categories. Notably, this effect was consistent regardless of whether or not the image included textual annotations ( for both cases).
@@
Our analysis shows that all variable groups contribute to higher perceived visual complexity: information-theoretic metrics, greater edge density, more feature points, and a larger number of distinct objects. This result replicates the findings in vision science that visual complexity is likely to be a high-dimensional phenomenon influenced by many factors. As shown in the three panels of What Makes a Visualization Image Complex?, the simplest and most complex visualizations differ notably in visual content. Similar patterns are observed in Apdx. Figure 12 when comparing samples from MASSVIS and VIS30K. Some visualizations, particularly matrix views, appear more complex due to their use of many small points and a wider range of colors. In contrast, simpler images tend to contain fewer elements and use minimal colors. Among these color-rich images, some used discrete colormaps with clearly defined edges, while others employed continuous colormaps such as gradient-style heatmaps. In some cases, these color patterns also reinforced coherent textural structures, further adding to visual complexity. Due to these observations, we next separate color/texture images by continuity to further investigate their relationship with perceived VC.